OpenStudy (anonymous):

prove the trig functions and show the steps 0=theta when i write this 6. sec0/tan0=csc0 7. (sec x +tan x)(1-sin x/cos x)=1 8. 2/sqrt3 cosx+sinx=sec(pi/6-x) 9. prove tan(0/2)=sin0/1+cos0 for 0 in quadrant 1 by filling in the reason belo

OpenStudy (mertsj):

\[\frac{\sec \theta}{\tan \theta}=\csc \theta\] \[\frac{1}{\cos \theta}\over\frac{\sin \theta}{\cos \theta}\] \[(\frac{1}{\cos \theta})\times \cos \theta \over(\frac{\sin \theta}{\cos \theta} )\times \cos \theta\]

OpenStudy (mertsj):

\[\frac{1}{\sin \theta}=\csc \theta\]

OpenStudy (mertsj):

\[(secx+tanx)(\frac{1-sinx}{cosx})=1\] \[(\frac{1}{cosx}+\frac{sinx}{cosx})(\frac{1-sinx}{cosx})\] \[\frac{1-sinx}{\cos ^{2}x}+\frac{sinx-\sin ^{2}x}{\cos ^{2}x}\] \[\frac{1-sinx+sinx-\sin ^{2}x}{\cos ^{2}x}\] \[\frac{1-\sin ^{2}x}{\cos ^{2}x}\] \[\frac{\cos^2x}{\cos^2x}=1\]

OpenStudy (mertsj):

Can't tell what you mean by the third one. Don't know if sinx is part of the denominator or not.